# Energy-and Spectral-Efficiency Trade-off in Distributed Massive-MIMO Networks

### v4

@@ I Introduction
Unlike centralized MIMO systems where transmit power dominates total energy consumption, D-mMIMO introduces significant additional energy costs due to distributed circuit operations, fronthaul communication, and cooperative signal processing [3]. In such systems, the total number of active APs and their associations with UEs have a direct and substantial influence on system-wide energy consumption. Therefore, evaluating the EE-SE trade-off in D-mMIMO systems from a system design perspective necessitates incorporating these architecture-specific energy components.
@@ IV Fractional Programming-Based Energy Efficiency Maximization
Step 1: Reformulating the Objective Function: In the original objective function (11a), both the numerator and the denominator contain the sum SE term. This increases the complexity of solving the problem. Therefore, without affecting the optimality of the solution, we reformulate the objective by removing the SE-dependent term from the denominator. Specifically, we consider a simplified energy efficiency expression where the denominator only includes the static and circuit power consumption terms. The modified optimization problem becomes:
@@
Since the removed term $P^{\text{deco}}_{cpu}B\sum_{t=1}^{T}\text{SE}^{u}_{t}$ is monotonically increasing with respect to the sum SE, maximizing the modified objective function in (12a) still leads to the same optimal solution as the original formulation in (11a). Therefore, solving Problem (12) is equivalent to solving the original Problem (11), but with a lower computational cost.
@@
: The quadratic transformation in (16) provides a concave lower bound for $\Gamma_{t}$. Therefore, constraint (15) serves as a approximation of the original non-convex constraint (14), and its substitution preserves the feasibility of Problem (12). Furthermore, when $z_{t}=z^{*}_{t}$, the lower bound becomes tight and constraint (15) is equivalent to (14), where
@@
Step 6: Iterative Solution: After solving Problem (24), the auxiliary variables $b$ and z are updated, and Problem (23) is solved again. This alternating procedure between the two sub-problems continues until convergence. The iterative procedure is summarized in Algorithm 1.
@@ V Numerical Simulations
We conduct numerical simulations over a $1\times 1$ km2 geographic area where APs and UEs are randomly distributed using the uniform distribution. To emulate an infinitely large network and eliminate edge effects, we adopt a wrap-around topology as described in [15]. In our setup, the number of APs is denoted by $M$, and the number of UEs is $T$.
The large-scale fading coefficients are modeled using a three-slope path loss model, with shadow fading applied using a log-normal distribution with 8 dB standard deviation. Unless specified otherwise, we fix the system parameters to $A=8$, $L_{c}=200$, and $L_{p}=5$. During the channel estimation phase, all pilots are transmitted at the maximum power of 100 mW. The initial UE transmit power for data transmission is also set to 100 mW. For the initial AP-UE association, each AP serves only those UEs for which it contributes at least 95% of the maximum large-scale fading coefficient (LSFC) across all APs.
Unless stated otherwise, all other parameters follow the configuration described in [1]. All numerical results are averaged over 50 independent simulation realizations.
@@ VI Conclusion and Future Work
From a system design perspective, the proposed framework provides essential guidance for deploying scalable and sustainable D-mMIMO networks. It enables network operators to adaptively activate or deactivate APs based on user density and SE demands, thereby avoiding energy over-provisioning and enhancing operational efficiency. As future wireless networks such as 5G-Advanced and 6G increasingly prioritize high throughput and low latency, the insights from this work may be useful in highlighting the critical importance of integrating EE considerations into the core of radio resource management and infrastructure planning. Ultimately, this work supports the design of green and economically viable wireless communication systems without compromising quality of service.
A compelling future direction would involve drawing detailed comparisons of the EE-SE trade-offs across different network architectures, including small-cell networks, traditional mMIMO and the D-mMIMO. Additionally, the influence of modulation and coding schemes, which is not considered in this study, presents an interesting direction for further exploration.

### v5

@@ I Introduction
Unlike centralized MIMO systems where transmit power dominates total energy consumption, D-mMIMO introduces significant additional energy costs due to distributed circuit operations, fronthaul communication, and cooperative signal processing [3]. In such systems, the total number of active APs and their associations with UEs have a more significant impact on system-wide energy consumption than the over-the-air transmit power. Therefore, evaluating the EE-SE trade-off in D-mMIMO systems from a system design perspective necessitates incorporating these architecture-specific energy components.
@@ IV Fractional Programming-Based Energy Efficiency Maximization
Step 1: Reformulating the Objective Function: In the original objective function (11a), both the numerator and the denominator contain the sum SE term. This increases the complexity of solving the problem. Therefore, without affecting the optimality of the solution, we reformulate the objective by removing the $P^{\text{deco}}_{cpu}wB\sum_{t=1}^{T}\log_{2}(1+\Gamma_{t})$ term, which is monotonically increasing with respect to the sum SE, from the denominator. Specifically, we consider a simplified energy efficiency expression where the denominator only includes the static and circuit power consumption terms. The modified optimization problem becomes:
@@
Since the removed term $P^{\text{deco}}_{cpu}wB\sum_{t=1}^{T}\log_{2}(1+\Gamma_{t})$ is monotonically increasing with respect to the sum SE, maximizing the modified objective function in (12a) still leads to the same optimal solution as the original formulation in (11a). Therefore, solving Problem (12) is equivalent to solving the original Problem (11), but with a lower computational cost.
@@
: The quadratic transformation in (16) provides a concave lower bound for $\Gamma_{t}$. Therefore, constraint (15) serves as an approximation of the original non-convex constraint (14), and its substitution preserves the feasibility of Problem (12). Furthermore, when $z_{t}=z^{*}_{t}$, the lower bound becomes tight and constraint (15) is equivalent to (14), where
@@
Step 6: Iterative Solution: After solving Problem (24), the auxiliary variables $b$ and z are updated, and Problem (23) is solved again. This alternating procedure between the two sub-problems continues until convergence. The complete iterative procedure is summarized in Algorithm 1, and their visual representation of steps involved for optimization is provided in Fig. 2.
@@
Start / System ParametersReformulate Objective Function :  $\frac{wB\sum_{t=1}^{T}\log_{2}(1+\Gamma_{t})}{P^{\text{fix}}_{\text{T}}+{P}^{c}_{\text{T}}+P^{\text{deco}}_{cpu}wB\sum_{t=1}^{T}\log_{2}(1+\Gamma_{t})}\xRightarrow{}\frac{wB\sum_{t=1}^{T}\log_{2}(1+\Gamma_{t})}{P^{\text{fix}}_{\text{T}}+{P}^{c}_{\text{T}}}$elax Binary AP-UE Variables : $d_{mk}\in\{0,1\}\xRightarrow{}d_{mk}\in[0,1]$Apply Quadratic Transformation for SINR : $\frac{\mathrm{DS}_{t}}{I_{t}}\xRightarrow{}2z_{t}\sqrt{\mathrm{DS}_{t}}-z_{t}^{2}I_{t}$gain Apply Quadratic Transformation for EE Fraction : $\frac{u}{v}\xRightarrow{}2b\sqrt{u}-b^{2}v$Decouple Power Factor ($\bm{\eta^{u}}$) and Association Variables ($\mathbf{D}$) by Solving Optimization AlternativelyIterative Refinement and Convergence CheckConverged EE
Figure 2: Flowchart of the proposed optimization framework for EE-SE trade-off in D-mMIMO systems.
@@ V Numerical Simulations
We conduct numerical simulations over a $1\times 1$ km2 geographic area where APs and UEs are independently and identically distributed over a square area, centred at the origin. Specifically, the x and y coordinates of each UE and AP are independently drawn from a continuous uniform distribution ranging from -0.5 to 0.5. To emulate an infinitely large network and eliminate edge effects, we adopt a wrap-around topology as described in [15]. In our setup, the number of APs is denoted by $M$, and the number of UEs is $T$.
The large-scale fading coefficients are modeled using a three-slope path loss model, with shadow fading applied using a log-normal distribution with 8 dB standard deviation. Unless specified otherwise, we fix the system parameters to $B=20$MHz $A=8$, $L_{c}=200$, and $L_{p}=5$. During the channel estimation phase, all pilots are transmitted at the maximum power of 100 mW. The initial UE transmit power for data transmission is also set to 100 mW. For the initial AP-UE association, each AP serves only those UEs for which it contributes at least 95% of the maximum large-scale fading coefficient (LSFC) across all APs. Also, $P^{\text{c}}_{ue}=100$ mW , $P^{\text{c}}_{ap}=100$ mW, $P^{\text{proc}}=800$, mW $P^{\text{fix}}_{fh}=825$ mW, $P^{\text{sig}}=10$ mW, $P^{\text{fix}}_{cpu}=5000$ mW, $P^{\text{lsfd}}_{cpu}=1000$, and $P^{\text{deco}}_{cpu}=1000$ mW/Gb/s [3]. Unless stated otherwise, all other parameters follow the configuration described in [1]. All numerical results are averaged over 50 independent simulation realizations.
@@ VI Conclusion and Future Work
From a system design perspective, the proposed framework provides essential guidance for deploying scalable and sustainable D-mMIMO networks. It enables network operators to adaptively activate or deactivate APs based on user density and SE demands, thereby avoiding energy over-provisioning and enhancing operational efficiency. Moreover, the AP-UE association plays a critical role in balancing fronthaul signaling power, further influencing overall energy consumption and coordination overhead. As future wireless networks such as 5G-Advanced and 6G increasingly prioritize high throughput and low latency, the insights from this work may be useful in highlighting the critical importance of integrating EE considerations into the core of radio resource management and infrastructure planning. Ultimately, this work supports the design of green and economically viable wireless communication systems without compromising quality of service.
A compelling future direction would involve drawing detailed comparisons of the EE-SE trade-offs across different network architectures, including small-cell networks, traditional mMIMO and the D-mMIMO. Additionally, the influence of modulation and coding schemes, which is not considered in this study, presents an interesting direction for further exploration. Furthermore, incorporating dynamic traffic patterns could provide deeper insights into real-time energy-efficient network control.